Back to main Archive News page.
Tuesday 31 May 2005 - Online Payments
The online payment service has now been restored and is once again available for paying bills or the topping up of print credit. Printing credit can also be topped up using cash or cheque at ISS Service Desk in the Library.
Monday 30 May 2005 - All systems ran without major problems
Friday 27 May 2005 - New Email Virus Claiming To Come From 'Admin@lancaster.ac.uk
The W32.MyDoom.BU@mm email virus is circulating on campus. It has various subject lines including: 'Your Email Account is Suspended For Security Reasons'. A typical body text is: 'Please read the attached document and follow it's instructions.'. Do NOT open the attachment in this email or any similiar messages. Any real messages about email accounts would come from the the ISS Service Desk. To read about W32.MyDoom.BU@mm visit this link.
Thursday 26 May 2005 - All systems ran without major problems
Wednesday 25 May 2005 - Online Payment Problems
There is currently a problem with the Online Payments system. This is affecting users wishing to pay bills online, or top-up their print credit. You can still top up your print credit by cash or cheque in person at ISS Reception in the Library.
Tuesday 24 May 2005 - All systems ran without major problems
Monday 23 May 2005 - Disconnected Computers
Following a security problem on campus on Friday, several office computers have had their network connections turned off. If you suspect your computer has been disconnected please contact the ISS Service Desk on x10987 for advice. Do not plug it into the next data point (and the one after that) as this could cause severe problems for other users on campus, and will certainly slow down the fixing process.
Friday 20 May 2005 -
Thursday 19 May 2005 - All systems ran without major problems
Wednesday 18 May 2005 - All systems ran without major problems
Tuesday 17 May 2005 - Network Problems in Grad College
Monday 16 May 2005 - New version of W32.MyDoom seen
The W32.MyDoom.BT@MM virus has a subject line of "FW: EMAIL ACCOUNT SUSPENSION"
(or simliar) and appears to come from "Admin@lancaster.ac.uk". Do
Friday 13 May 2005 - All systems ran without major problems.
Thursday 12 May 2005 - All systems ran without major problems
Wednesday 11 May 2005 - All systems ran without major problems
Tuesday 10 May 2005 - Network Problems in Grad College
Over the weekend a media converter failed in Southwest campus. This was replaced on Sunday afternoon, however the replacement (media converter) failed later on the Sunday. Reports of problems were made to ISS on Tuesday at 16:45. A new solution was found and service was restored properly at 19:00 that same evening. Reliability is now much improved
Monday 09 May 2005 - All systems ran without major problems
Friday 06 May 2005 - All systems ran without major problems
Thursday 05 May 2005 - Excessive network traffic caused the network to slow
There was excessive network traffic ('broadcast storm') causing slow response times and/or difficulty in accessing network resources including MS-Exchange, H: Drive, etc. A configuration change has been made on the core switches which has resolved the problem. To further investigate the problem, scheduled maintenance will be undertaken during the next network at risk period.
Wednesday 04 May 2005 - All systems ran without major problems
Tuesday 03 May 2005 - Network problems requiring core switch change
The regular network at risk period saw one of the core network switches replaced. Due to the complexity of the job, completion was scheduled for 10am, but was actually finished by 09:15. As expected, the work did disrupt network connectivity across campus and to the wider internet whilst being carried out.
The new core switch replaced this morning has experienced intermittent problems, causing users to be unable to connect to various network resources (MS-Exchange, H: drive, etc). As at 15:15 there were no outstanding problems. The outage appears to have been caused by the new core switch failing to update its forwarding database correctly. Following a reboot this problem has not been seen again. Having an incorrect and incomplete forwarding database will cause what first appears to be a broadcast storm.
Monday 02 May 2005 - All systems ran without major problems